OpenStudy (anonymous):

Factor completely c^2+9-6c

OpenStudy (anonymous):

rearrange as c^2 - 6c + 9 = c^2 - 3c -3c + 9 = c(c - 3) -3(c - 3) = (c-3)(c-3)